OpenStudy (anonymous):

A contractor employed 210 men to build a house in 60 days. After 12 days , he was joined by 70 more men. In how many days will the remaining work be finished.

OpenStudy (anonymous):

Come on, 1/3 more men to finish 4/5 of work....

OpenStudy (anonymous):

How many mandays the job takes?

OpenStudy (anonymous):

if 210 men take 60 days then 280 men would take 210/280 * 60 = 45 days that is 280 take 3/4 of the time that 210 take and there are 48 days left if 210 do the work so how do you calculate how long 280 men would take?

OpenStudy (amistre64):

the standard reply to any contracting job is: itll be finished in 2 weeks. :)

OpenStudy (anonymous):

And at 10 days, u know, there's a bit of a problem.........

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

her each man cando 1/60*280 part of job per day so 210 of them worked for 12 days and completed 12*210/60*210 =1/5 th part of the job now job ramaining is 4/5th they have 70men more and total of 280 let the number of days rquired be x then x*280/60*210=4/5 so x=180/5=36 so they take 36 more days to finish the work